OpenStudy (anonymous):

in isosceles PQR, P is the vertex angle. If M Q =8x-3 and mR = 2x+15, find the measure of P

OpenStudy (phi):

I assume mQ means measure of angle Q the base angles of an isosceles triangle are equal. that means 8x-3 = 2x+15 can you "solve for x"?

OpenStudy (anonymous):

thanks and yes i can

OpenStudy (phi):

this is multiple steps. once you find x, you have to find mQ (which also be mR) then find the 3rd angle, which is what they want.

OpenStudy (anonymous):

i dont get that part.

OpenStudy (phi):

did you find a value for x ?

OpenStudy (anonymous):

yes

OpenStudy (phi):

and x = ?

OpenStudy (anonymous):

2

OpenStudy (phi):

try again

OpenStudy (anonymous):

its wrong.?

OpenStudy (phi):

if x =2 then 8x-3 = 2x+15 becomes 8*2 - 3 = 2*2 +15 16-3 = 4+15 13 = 19 , which is not true.

OpenStudy (anonymous):

oh thanks i get it now